You are the manager of a monopoly firm with (inverse) demand given by P = 50 − 0.5Q. Your firm's cost function is C = 40 + 5Q2. Your firm's marginal revenue is:

Multiple Choice

P = 50 − 0.5Q.

P = 50 − Q.

There is insufficient information to determine the firm's marginal revenue.

P = 100 − Q.


1
Expert's answer
2021-06-21T12:20:46-0400
"TR=pQ=(50-0.5Q)Q=50Q-0.5Q^2"

"MR=\frac{\varDelta TR}{\varDelta Q}=50-0.5\times2Q=50-Q"